OpenStudy (alexwee123):

a spiiner has 40% of it colored blue. Red and green make up the rest of the spinner in equal amounts. A) if the spinner is spun twice, what is the probability of getting the same color on both spins? b) suppose you know the spinner landed on the same color twice. What is the probability that the color was red? c) Suppose you know that at least one time the spinner landed on blue when it was spun twice. What is the probability that the other color was also blue?

OpenStudy (anonymous):

is it to use this equation y=e-^tk

OpenStudy (alexwee123):

nope this is probability :(

OpenStudy (anonymous):

oh i though calcus

OpenStudy (anonymous):

for A) if the spinner is spun twice, what is the probability of getting the same color on both spins? There are 3 situations that can occur, both spins are blue, both spins are green or both spins are red. Probability = P(B,B) + P(G,G) + P(R,R) = 0.4*0.4 + 0.3*0.3 + 0.3*0.3 = 0.34 b) suppose you know the spinner landed on the same color twice. What is the probability that the color was red? P( Red, given that spinner landed same color twice) = P(Both spins are red)/ P(Both spins are same color, the ans in part A) = 0.3*0.3 / 0.34 = 0.26 c) Suppose you know that at least one time the spinner landed on blue when it was spun twice. What is the probability that the other color was also blue? P ( both blue, given that at least one was blue) = P(B,B) / [P(B,B) + P(B,any color)] = 0.16/ [0.4*0.4 + 0.4*0.6] = 0.38

OpenStudy (alexwee123):

oh i get it it :D thx